# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@
We start with some basic notation that we use throughout the whole paper.
@@
    For all $\mathbb{K}\in\{\mathbb{N},\mathbb{R}\}$ and $a\in\mathbb{R}$ set $\mathbb{K}_{\geq a}:=\mathbb{K}\cap[a,\infty)$ and $\mathbb{K}_{>a}:=\mathbb{K}\cap(a,\infty)$. Analogously, $\mathbb{K}_{\leq a}$ and $\mathbb{K}_{<a}$ are defined. In particular, we set $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
@@
    For all $a,b\in\mathbb{Z}$ with $a\leq b$ set $\left\llbracket a,b\right\rrbracket:=[a,b]\cap\mathbb{Z}$ and further $\left\llbracket n\right\rrbracket:=\left\llbracket 1,n\right\rrbracket$ for any $n\in\mathbb{N}$.
@@
    For two normed spaces $(S_{1},\left\|\cdot\right\|_{S_{1}})$ and $(S_{2},\left\|\cdot\right\|_{S_{2}})$, we write $S_{1}\hookrightarrow S_{2}$, if $S_{1}\subseteq S_{2}$ is continuously embedded into $S_{2}$.
@@
    For $d\in\mathbb{N}$ denote by $\left\|\cdot\right\|_{\mathbb{R}^{d}}$ the usual Euclidean norm on $\mathbb{R}^{d}$.
@@
    For a function $f:\mathcal{S}\to\mathbb{R}$ defined on an arbitrary space $\mathcal{S}$, we denote by $\left\|f\right\|_{\infty}:=\sup_{s\in\mathcal{S}}\left\{|f(s)|\right\}$ the supremum norm of $f$ over $\mathcal{S}$. We do not highlight the underlying space, if it is clear from the context.
  8. viii)
@@
   Hölder and Sobolev spaces
Throughout let $(\mathcal{Z},\mathscr{Z},\zeta)$ be a measure space. For $p\in[1,\infty]$ we define the usual space $\mathbb{L}_{\zeta}^{p}(\mathcal{Z}):=\mathbb{L}^{p}(\mathcal{Z},\mathscr{Z},\zeta)$ of (equivalence classes of) functions $f:(\mathcal{Z},\mathscr{Z})\to(\mathbb{R},\mathscr{B}_{\mathbb{R}})$ endowed with the norm $\left\|\cdot\right\|_{\mathbb{L}^{p}_{\zeta}(\mathcal{Z})}$. Let $(V,\left|\cdot\right|_{V})$ be a finite-dimensional normed $\mathbb{R}$-vector space with dimension $d_{V}:=\operatorname{dim}_{\mathbb{R}}(V)$. By identifying $V$ canonically with $\mathbb{R}^{d_{V}}$, we define the space of all vector fields $f=(f_{1},\dots,f_{d_{V}}):\mathcal{Z}\to V$ with the integrable components $f_{i}\in\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$, $i\in\left\llbracket d_{V}\right\rrbracket$, as the tensor space $\mathbb{L}^{p}_{\zeta}(\mathcal{Z},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$. Recalling that $\mathbb{L}^{2}_{\zeta}(\mathcal{Z})$ defines a Hilbert space with inner product
@@
where $\cdot_{i}$ indicates the $i$-th component. Throughout let $d\in\mathbb{N}$. We denote by $\mathcal{O}\subseteq\mathbb{R}^{d}$ a bounded domain with smooth boundary $\partial\mathcal{O}$. For any $s\in\mathbb{N}$ denote by $C^{s}(\mathcal{O})$ the space of $s$-times continuously differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with bounded derivatives. By convention, we denote by $C^{0}(\mathcal{O})$ the space of continuous functions on $\mathcal{O}$. We extend the definition for any $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by defining $C^{s}(\mathcal{O})$ as the function space of all maps $f:\mathcal{O}\to\mathbb{R}$, such that for all multi-indices $\bm{\beta}\in\mathbb{N}_{0}^{d}$ with $|\bm{\beta}|\leq\lfloor s\rfloor$ the integer part of $s$ the derivative $D^{\bm{\beta}}f$ exists and is $s-\lfloor s\rfloor$-Hölder continuous. A norm on $C^{s}(\mathcal{O})$ is given by
@@
$\left\|f\right\|_{C^{s}(\mathcal{O})}:=\sum_{\bm{\beta}\in\mathbb{N}_{0}^{d}:|\bm{\beta}|\leq\lfloor s\rfloor}\left\|D^{\bm{\beta}}f\right\|_{\infty}+\sum_{\bm{\beta}\in\mathbb{N}_{0}^{d}:|\bm{\beta}|=\lfloor s\rfloor}\sup_{x,y\in\mathcal{O}:x\neq y}\left\{\frac{|D^{\bm{\beta}}f(x)-D^{\bm{\beta}}f(y)|}{\left\|x-y\right\|_{\mathbb{R}^{d}}^{s-\lfloor s\rfloor}}\right\}.$
@@
A function $f:\mathcal{O}\to\mathbb{R}$ is called smooth, if it belongs to $C^{\infty}(\mathcal{O}):=\bigcap_{s\in\mathbb{R}_{>0}}C^{s}(\mathcal{O})$. We further denote by $C^{\infty}_{c}(\mathcal{O})$ its subspace of all smooth functions having compact support. Further let $\mathcal{L}^{d}$ be the usual $d$-dimensional Lebesgue measure on $(\mathbb{R}^{d},\mathscr{B}_{\mathbb{R}}^{\otimes d})$ and denote by $\mathcal{L}_{\mathcal{O}}^{d}$ its restriction on $\mathcal{O}$. To shorten notation, set $\mathbb{L}^{2}(\mathcal{O}):=\mathbb{L}^{2}_{\mathcal{L}_{\mathcal{O}}^{d}}(\mathcal{O})$ and analogously $\mathbb{L}^{2}(\mathcal{O},V):=\mathbb{L}^{2}_{\mathcal{L}_{\mathcal{O}}^{d}}(\mathcal{O},V)$. For $s\in\mathbb{N}$, we denote by $H^{s}(\mathcal{O}):=H^{s}(\mathcal{O},\mathbb{R})$ the usual order-$s$ Sobolev space of weakly differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with square integrable weak derivatives up to order $s$. By $\left\|\cdot\right\|_{H^{s}(\mathcal{O})}$, we denote its usual norm. This definition extends to positive $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by interpolation, see Triebel [1983]. By convention, set $H^{0}(\mathcal{O}):=\mathbb{L}^{2}(\mathcal{O})$. Further define $H^{s}_{c}(\mathcal{O}):=\overline{C_{c}^{\infty}(\mathcal{O})}^{\left\|\cdot\right\|_{H^{s}(\mathcal{O})}}$ as well as $\bar{H}^{s}(\mathcal{O})$ to be the space of all functions in $H^{s}(\mathcal{O})$, whose support is $\bar{\mathcal{O}}$. If $s\neq k+\frac{1}{2}$ for some $k\in\mathbb{N}$, we have $\bar{H}^{s}(\mathcal{O})=H^{s}_{c}(\mathcal{O})$. Negative-order Sobolev spaces are defined as the topological dual of $\bar{H}^{s}(\mathcal{O})$, i.e., $H^{-s}(\mathcal{O}):=H^{-s}(\mathcal{O},\mathbb{R}):=\left(\bar{H}^{s}(\mathcal{O})\right)^{*}$. Denoting by $\operatorname{tr}_{s}:H^{s}(\mathcal{O})\to\bigtimes_{l=0}^{s-1}H^{s-l-\frac{1}{2}}(\mathcal{O})$ the higher-order trace operator for any $s\in\mathbb{N}$, we have $H^{s}_{0}(\mathcal{O}):=\left\{f\in H^{s}(\mathcal{O}):\,\operatorname{tr}_{s}[f]=0\right\}=H^{s}_{c}(\mathcal{O})$, where we refer to Lions and Magenes [1972] for detailed definitions and properties. All of the previous spaces of functions defined on $\mathcal{O}$ with values in $\mathbb{R}$ can be extended to spaces of the corresponding vector fields $f:\mathcal{O}\to V$ in the same manner as before. Indeed if $F(\mathcal{O})$ is one of the above function spaces, we set $F(\mathcal{O},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}F(\mathcal{O})$. If $F(\mathcal{O})$ is equipped with some norm $\left\|\cdot\right\|_{F(\mathcal{O})}$, a norm on $F(\mathcal{O},V)$ is given by $\left\|\cdot\right\|_{F(\mathcal{O},V)}^{2}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\left\|\cdot_{i}\right\|_{F(\mathcal{O})}^{2}$. Similarly, if $F(\mathcal{O})$ is endowed with a Hilbert space inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O})}$, $F(\mathcal{O},V)$ inherits the Hilbert space structure with the inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O},V)}:=\sum_{i\in\left\llbracket d_{V}\right\rrbracket}\langle\cdot_{i},\cdot_{i}\rangle_{F(\mathcal{O})}$. Further, for simplicity, we write for centered balls with radius $r\in\mathbb{R}_{>0}$ in function spaces $F(\mathcal{O})$ and $F(\mathcal{O},V)$ $F(\mathcal{O},r):=\mathbb{B}_{F(\mathcal{O})}(0,r)$ and $F(\mathcal{O},V,r):=\mathbb{B}_{F(\mathcal{O},V)}(0,r)$, respectively.
   The measurement model
@@
Throughout let $(W,\left|\cdot\right|_{W})$ be another finite-dimensional and normed $\mathbb{R}$-vector space with dimension $d_{W}:=\operatorname{dim}_{\mathbb{R}}(W)\in\mathbb{N}$, such that we canonically identify $W$ with $\mathbb{R}^{d_{W}}$. Assuming that $(\mathcal{Z},\mathscr{Z},\zeta)$ is a probability space, we are interested in a (possibly nonlinear) forward map
@@
$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i\in\left\llbracket N\right\rrbracket,\,\theta\in\tilde{\Theta},$ (2.1)
where $Z_{1},\dots,Z_{N}$ are i.i.d copies of $Z$, which is drawn from $\zeta$. The error terms $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d. copies of $\varepsilon\sim\operatorname{N}(0,\operatorname{Id}_{V})$ with $\operatorname{Id}_{V}$ denoting the identity operator on the vector space $V$. The design variable $Z$ and the error $\varepsilon$ are assumed to be independent. The distribution of a generic observation $D:=(Y,Z)$ arising from Eq. 2.1 is denoted by $\mathbb{P}_{\theta}$. The joint product law of $D_{N}$ on $(V\times\mathcal{Z})^{N}$ is denoted as $\mathbb{P}_{\theta}^{N}=\bigotimes_{i\in\left\llbracket N\right\rrbracket}\mathbb{P}_{\theta}$. The corresponding expectation operators are denoted by $\mathbb{E}_{\theta}$ and $\mathbb{E}_{\theta}^{N}$, respectively. Keeping the identification of $V$ with $\mathbb{R}^{d_{V}}$ in mind, $\mathbb{P}_{\theta}$ has a probability density function $p_{\theta}$ with respect to the product measure $\mathcal{L}^{d_{V}}\otimes\zeta$, namely
$\forall y\in V,z\in\mathcal{Z}:\,p_{\theta}(y,z)=(2\pi\sigma^{2})^{-\frac{d_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\left|y-\mathscr{G}(\theta)(z)\right|_{V}^{2}\right).$
@@
$p_{\theta}^{N}(y,z)=\prod_{i\in\left\llbracket N\right\rrbracket}p_{\theta}(y_{i},z_{i})=(2\pi\sigma^{2})^{-\frac{Nd_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|y_{i}-\mathscr{G}(\theta)(z_{i})\right|_{V}^{2}\right).$
@@
$\ell_{N}(\theta):=\ell_{N}(\theta|D_{N}):=\log p_{\theta}^{N}(D_{N})+\frac{Nd_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}$
@@
$\ell(\theta):=\ell(\theta|D):=\log p_{\theta}(D)+\frac{d_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\left|Y-\mathscr{G}(\theta)(Z)\right|_{V}^{2}.$
The results obtained below hold under the frequentist assumption that the data $D_{N}$ from Eq. 2.1 is generated from a fixed ground truth $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ including the law $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$.
   Maximum a posteriori estimator
We are now introducing the maximum a posteriori estimator, whose statistical guarantees are presented in the following sections. Given the data set $D_{N}=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ arising from Eq. 2.1, fix some $\alpha,\mathfrak{r}\in\mathbb{R}_{\geq 0}$. We define a penalized least squares functional by
$\mathscr{J}_{\mathfrak{r},N}:\tilde{\Theta}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\text{ }\mathscr{J}_{\mathfrak{r},N}[\theta]:=-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}\left|Y_{i}-\mathscr{G}(\theta)(Z_{i})\right|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\theta\right\|_{H^{\alpha}(\mathcal{O},W)}^{2}$ (2.2)

### v3

@@ 2 A general statistical inverse Problem
Throughout the whole paper, we use the following notation.
@@
    For $\mathbb{K}\in\{\mathbb{N},\mathbb{R}\}$ and $a\in\mathbb{R}$ set $\mathbb{K}_{\geq a}:=\mathbb{K}\cap[a,\infty)$ and $\mathbb{K}_{>a}:=\mathbb{K}\cap(a,\infty)$. Analogously, $\mathbb{K}_{\leq a}$ and $\mathbb{K}_{<a}$ are defined. We set $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$.
@@
    For all $a,b\in\mathbb{Z}$ with $a\leq b$ set $\left\llbracket a,b\right\rrbracket:=[a,b]\cap\mathbb{Z}$, and $\left\llbracket n\right\rrbracket:=\left\llbracket 1,n\right\rrbracket$ for any $n\in\mathbb{N}$.
@@
    For two normed spaces $(S_{1},\|\cdot\|_{S_{1}})$ and $(S_{2},\|\cdot\|_{S_{2}})$, we write $S_{1}\hookrightarrow S_{2}$, if $S_{1}$ is continuously embedded into $S_{2}$.
@@
    For $d\in\mathbb{N}$ denote by $\|\cdot\|_{\mathbb{R}^{d}}$ the usual Euclidean norm on $\mathbb{R}^{d}$.
@@
2.1 Hölder and Sobolev spaces
Throughout let $(\mathcal{Z},\mathscr{Z},\zeta)$ be a measure space. For $p\in[1,\infty]$ we define the usual space $\mathbb{L}_{\zeta}^{p}(\mathcal{Z}):=\mathbb{L}^{p}(\mathcal{Z},\mathscr{Z},\zeta)$ of (equivalence classes of) functions $f:(\mathcal{Z},\mathscr{Z})\to(\mathbb{R},\mathscr{B}_{\mathbb{R}})$ endowed with the norm $\|\cdot\|_{\mathbb{L}^{p}_{\zeta}(\mathcal{Z})}$. Let $(V,|\cdot|_{V})$ be a finite-dimensional normed $\mathbb{R}$-vector space with dimension $d_{V}:=\operatorname{dim}_{\mathbb{R}}(V)$. By identifying $V$ canonically with $\mathbb{R}^{d_{V}}$, we define the space of all vector fields $f=(f_{1},\dots,f_{d_{V}}):\mathcal{Z}\to V$ with the integrable components $f_{i}\in\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$, $i\in\left\llbracket d_{V}\right\rrbracket$, as the tensor space $\mathbb{L}^{p}_{\zeta}(\mathcal{Z},V):=\bigtimes_{i\in\left\llbracket d_{V}\right\rrbracket}\mathbb{L}^{p}_{\zeta}(\mathcal{Z})$. Recalling that $\mathbb{L}^{2}_{\zeta}(\mathcal{Z})$ defines a Hilbert space with inner product
@@ 2.1 Hölder and Sobolev spaces
where $\cdot_{i}$ indicates the $i$-th component. Moreover, for any function $f:\mathcal{Z}\to V$, we denote by $\|f\|_{\infty}:=\sup_{z\in\mathcal{Z}}|f(z)|_{V}$ the uniform norm of $f$ over $\mathcal{Z}$. We do not highlight the underlying space, if it is clear from the context. Throughout let $d\in\mathbb{N}$. We denote by $\mathcal{O}\subseteq\mathbb{R}^{d}$ a bounded domain with smooth boundary $\partial\mathcal{O}$. For any $s\in\mathbb{N}$ denote by $C^{s}(\mathcal{O})$ the space of $s$-times continuously differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with bounded derivatives. By convention, we denote by $C^{0}(\mathcal{O})$ the space of continuous functions on $\mathcal{O}$. We extend the definition for any $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by defining $C^{s}(\mathcal{O})$ as the function space of all maps $f:\mathcal{O}\to\mathbb{R}$, such that for all multi-indices $\boldsymbol{\beta}\in\mathbb{N}_{0}^{d}$ with $|\boldsymbol{\beta}|\leq\lfloor s\rfloor$ - the integer part of $s$ - the derivative $D^{\boldsymbol{\beta}}f$ exists and is $s-\lfloor s\rfloor$-Hölder continuous. A norm on $C^{s}(\mathcal{O})$ is given by
@@
$\|f\|_{C^{s}(\mathcal{O})}:=\sum_{\boldsymbol{\beta}\in\mathbb{N}_{0}^{d}:|\boldsymbol{\beta}|\leq\lfloor s\rfloor}\|D^{\boldsymbol{\beta}}f\|_{\infty}+\sum_{\boldsymbol{\beta}\in\mathbb{N}_{0}^{d}:|\boldsymbol{\beta}|=\lfloor s\rfloor}\sup_{x,y\in\mathcal{O}:x\neq y}\left\{\frac{|D^{\boldsymbol{\beta}}f(x)-D^{\boldsymbol{\beta}}f(y)|}{\|x-y\|_{\mathbb{R}^{d}}^{s-\lfloor s\rfloor}}\right\}.$
@@
A function $f:\mathcal{O}\to\mathbb{R}$ is called smooth, if it belongs to $C^{\infty}(\mathcal{O}):=\bigcap_{s\in\mathbb{R}_{>0}}C^{s}(\mathcal{O})$. We further denote by $C^{\infty}_{c}(\mathcal{O})$ its subspace of all smooth functions having compact support. Further let $\mathcal{L}^{d}$ be the usual $d$-dimensional Lebesgue measure on $(\mathbb{R}^{d},\mathscr{B}_{\mathbb{R}}^{\otimes d})$ and denote by $\mathcal{L}_{\mathcal{O}}^{d}$ its restriction on $\mathcal{O}$. To shorten notation, set $\mathbb{L}^{2}(\mathcal{O}):=\mathbb{L}^{2}_{\mathcal{L}_{\mathcal{O}}^{d}}(\mathcal{O})$ and analogously $\mathbb{L}^{2}(\mathcal{O},V):=\mathbb{L}^{2}_{\mathcal{L}_{\mathcal{O}}^{d}}(\mathcal{O},V)$. For $s\in\mathbb{N}$, we denote by $H^{s}(\mathcal{O}):=H^{s}(\mathcal{O},\mathbb{R})$ the usual order-$s$ Sobolev space of weakly differentiable functions $f:\mathcal{O}\to\mathbb{R}$ with square integrable weak derivatives up to order $s$. By $\|\cdot\|_{H^{s}(\mathcal{O})}$, we denote its usual norm. This definition extends to positive $s\in\mathbb{R}_{>0}\setminus\mathbb{N}$ by interpolation, see Triebel (1983). By convention, set $H^{0}(\mathcal{O}):=\mathbb{L}^{2}(\mathcal{O})$. Further define $H^{s}_{c}(\mathcal{O}):=\overline{C_{c}^{\infty}(\mathcal{O})}^{\|\cdot\|_{H^{s}(\mathcal{O})}}$ as well as $\bar{H}^{s}(\mathcal{O})$ to be the space of all functions in $H^{s}(\mathcal{O})$, whose support is $\bar{\mathcal{O}}$. If $s\neq k+\frac{1}{2}$ for any $k\in\mathbb{N}$, we have $\bar{H}^{s}(\mathcal{O})=H^{s}_{c}(\mathcal{O})$. Negative-order Sobolev spaces are defined as the topological dual of $\bar{H}^{s}(\mathcal{O})$, i.e., $H^{-s}(\mathcal{O}):=H^{-s}(\mathcal{O},\mathbb{R}):=\left(\bar{H}^{s}(\mathcal{O})\right)^{*}$. Denoting by $\operatorname{tr}_{s}:H^{s}(\mathcal{O})\to\bigtimes_{l=0}^{s-1}H^{s-l-\frac{1}{2}}(\mathcal{O})$ the higher-order trace operator for any $s\in\mathbb{N}$, we have $H^{s}_{0}(\mathcal{O}):=\left\{f\in H^{s}(\mathcal{O}):\,\operatorname{tr}_{s}[f]=0\right\}=H^{s}_{c}(\mathcal{O})$, where we refer to Lions and Magenes (1972) for detailed definitions and properties. Let $(W,\left|\cdot\right|_{W})$ be another finite-dimensional and normed $\mathbb{R}$-vector space with dimension $d_{W}:=\operatorname{dim}_{\mathbb{R}}(W)\in\mathbb{N}$, such that we canonically identify $W$ with $\mathbb{R}^{d_{W}}$. All of the previous spaces of functions defined on $\mathcal{O}$ with values in $\mathbb{R}$ can be extended to spaces of the corresponding vector fields $f:\mathcal{O}\to W$ in the same manner as before. Indeed, if $F(\mathcal{O})$ is one of the above function spaces, we set $F(\mathcal{O},W):=\bigtimes_{i\in\left\llbracket d_{W}\right\rrbracket}F(\mathcal{O})$. If $F(\mathcal{O})$ is equipped with some norm $\|\cdot\|_{F(\mathcal{O})}$, a norm on $F(\mathcal{O},W)$ is given by $\|\cdot\|_{F(\mathcal{O},W)}^{2}:=\sum_{i\in\left\llbracket d_{W}\right\rrbracket}\|\cdot_{i}\|_{F(\mathcal{O})}^{2}$. Similarly, if $F(\mathcal{O})$ is endowed with a Hilbert space inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O})}$, $F(\mathcal{O},W)$ inherits the Hilbert space structure with the inner product $\langle\cdot,\cdot\rangle_{F(\mathcal{O},W)}:=\sum_{i\in\left\llbracket d_{W}\right\rrbracket}\langle\cdot_{i},\cdot_{i}\rangle_{F(\mathcal{O})}$. Further, for simplicity, we write for centered balls with radius $r\in\mathbb{R}_{>0}$ in function spaces $F(\mathcal{O})$ and $F(\mathcal{O},W)$, $F(\mathcal{O},r):=\mathbb{B}_{F(\mathcal{O})}(0,r)$ and $F(\mathcal{O},W,r):=\mathbb{B}_{F(\mathcal{O},W)}(0,r)$, respectively.
2.2 The measurement model
@@
Assuming that $(\mathcal{Z},\mathscr{Z},\zeta)$ is a probability space, we are interested in a (possibly nonlinear) forward map
@@
$Y_{i}=\mathscr{G}(\theta)(Z_{i})+\sigma\varepsilon_{i},\,i\in\left\llbracket N\right\rrbracket,\,\theta\in\tilde{\Theta}$, (7)
where $Z_{1},\dots,Z_{N}$ are i.i.d copies of $Z$, which is drawn from $\zeta$. The error terms $\varepsilon_{1},\dots,\varepsilon_{N}$ are i.i.d. copies of $\varepsilon\sim\operatorname{N}(0,\operatorname{Id}_{V})$ with $\operatorname{Id}_{V}$ denoting the identity operator on the vector space $V$. The design variable $Z$ and the error $\varepsilon$ are assumed to be independent. Note, we assume that the noise variance $\sigma^{2}$ is known in advance. In settings where $\sigma^{2}$ is unknown, it could be estimated from the data, for instance via the sample variance. A detailed treatment of a unknown variance setting is beyond the scope of the present work and posponed to future work. The distribution of a generic observation $D:=(Y,Z)$ arising from eq. 7 is denoted by $\mathbb{P}_{\theta}$. The joint product law of $D_{N}$ on $(V\times\mathcal{Z})^{N}$ is denoted as $\mathbb{P}_{\theta}^{N}=\bigotimes_{i\in\left\llbracket N\right\rrbracket}\mathbb{P}_{\theta}$. The corresponding expectation operators are denoted by $\mathbb{E}_{\theta}$ and $\mathbb{E}_{\theta}^{N}$, respectively. Keeping the identification of $V$ with $\mathbb{R}^{d_{V}}$ in mind, $\mathbb{P}_{\theta}$ has a probability density function $p_{\theta}$ with respect to the product measure $\mathcal{L}^{d_{V}}\otimes\zeta$, namely
$\forall y\in V,z\in\mathcal{Z}:\,p_{\theta}(y,z)=(2\pi\sigma^{2})^{-\frac{d_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}|y-\mathscr{G}(\theta)(z)|_{V}^{2}\right).$
@@
$p_{\theta}^{N}(y,z)=\prod_{i\in\left\llbracket N\right\rrbracket}p_{\theta}(y_{i},z_{i})=(2\pi\sigma^{2})^{-\frac{Nd_{V}}{2}}\exp\left(-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|y_{i}-\mathscr{G}(\theta)(z_{i})|_{V}^{2}\right).$
@@
$\ell_{N}(\theta):=\ell_{N}(\theta|D_{N}):=\log p_{\theta}^{N}(D_{N})+\frac{Nd_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}$
@@
$\ell(\theta):=\ell(\theta|D):=\log p_{\theta}(D)+\frac{d_{V}}{2}\log(2\pi\sigma^{2})=-\frac{1}{2\sigma^{2}}|Y-\mathscr{G}(\theta)(Z)|_{V}^{2}.$
The results obtained below hold under the frequentist assumption that the data $D_{N}$ from eq. 7 is generated from a fixed ground truth $\theta_{\mathfrak{o}}\in\tilde{\Theta}$ including the law $\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}$.
2.3 Maximum a posteriori estimator
Given the data set $D_{N}=(Y_{i},Z_{i})_{i\in\left\llbracket N\right\rrbracket}\in(V\times\mathcal{Z})^{N}$ arising from eq. 7, fix some $\alpha,\mathfrak{r}\in\mathbb{R}_{\geq 0}$. We define a penalized least squares functional by
$\mathscr{J}_{\mathfrak{r},N}:\tilde{\Theta}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\text{ }\mathscr{J}_{\mathfrak{r},N}[\theta]:=-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-\mathscr{G}(\theta)(Z_{i})|_{V}^{2}-\frac{\mathfrak{r}^{2}}{2}\|\theta\|_{H^{\alpha}(\mathcal{O},W)}^{2}$ (8)